OpenStudy (anonymous):

–9 + (–5)x + 3 + (–2)x can u explain it plz

OpenStudy (anonymous):

consider the numbers in parenthesis as coefficients. then just combine like terms.

jigglypuff314 (jigglypuff314):

it can be looked at this way, -5x + (-2x) + (-9) + 3 ^ that part would be (-5 - 2)x = and then -9 + 3 =

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

so its –6 + 3x

jigglypuff314 (jigglypuff314):

-5 -2 is like -(5+2) =

OpenStudy (anonymous):

these are the options they gave me though A. –6 + 3x B. 3 – 16x C. –12 – 7x D. –6 – 7x

jigglypuff314 (jigglypuff314):

(-5 -2)x is like -(5+2)x = ___

OpenStudy (anonymous):

b

jigglypuff314 (jigglypuff314):

not quite... 5+2 = 7 right? so -(5+2)x = -7x so you get that -6 -7x

OpenStudy (anonymous):

ooh ty

jigglypuff314 (jigglypuff314):

you're welcome :)